### v1

# Divergence-Based Similarity Function for Multi-View Contrastive Learning

Abstract
Recent success in contrastive learning has sparked growing interest in more effectively leveraging multiple augmented views of an instance. While prior methods incorporate multiple views at the loss or feature level, they primarily capture pairwise relationships and fail to model the joint structure across all views. In this work, we propose a divergence-based similarity function (DSF) that explicitly captures the joint structure by representing each set of augmented views as a distribution and measuring similarity as the divergence between distributions. Extensive experiments demonstrate that DSF consistently improves performance across various tasks, including kNN classification and linear evaluation, while also offering greater efficiency compared to other multi-view methods. Furthermore, we establish a theoretical connection between DSF and cosine similarity, and show that, unlike cosine similarity, DSF operates effectively without requiring a temperature hyperparameter.
1 Introduction
Figure 1: Illustration of feature‐distribution modeling on the unit hypersphere: given multiple augmented views of the image, their extracted feature vectors (dots) lie on the hyperspherical manifold and are used to construct smooth probability distributions (colored regions) over the sphere.
Learning effective representations without costly labels is a central challenge in modern machine learning. Self-supervised learning (SSL) addresses this by deriving supervisory signals from the data itself, and contrastive learning has emerged as one of its most successful paradigms. In contrastive frameworks such as SimCLR[Chen et al., 2020a], MoCo[He et al., 2020], and DCL[Yeh et al., 2022], representations of two augmented views of the same instance are ”pulled together” while being ”pushed apart” from all other samples. A key ingredient across these methods is cosine similarity on the unit hypersphere, which normalizes features to lie on a sphere and then measures their dot-product similarity.
Figure 2: Illustration of multi-view contrastive learning strategies. A data instance is augmented $M$ times, resulting in $x^{1},\cdots,x^{M}$. Base encoder outputs $M$ representations $z^{1},\cdots,z^{M}$ from $M$ views. From top to bottom: (i) Loss_avg, which averages pairwise loss terms across $M$ views; (ii) Fea_avg, which computes a mean embedding over the $M$ views before applying contrast; and (iii) DSF, our proposed model that aggregates the view’s feature into distribution and fuses these distributions into a single contrastive loss
However, most existing contrastive methods focus on pairwise relationships between two views. Recent efforts to leverage more than two views can be broadly categorized into two approaches.:
  • Loss-level approaches: methods that formulate the contrastive objective by averaging pairwise losses computed between all available combinations of views, typically using arithmetic or geometric averaging schemes, as in [Shidani et al., 2024].
  • Feature-level approaches: methods that first aggregate multi-view representations—most notably through feature averaging—before applying a contrastive objective. For example, FastSiam [Pototzky et al., 2022] computes the average of view features, while other techniques like M3G[Piran et al., 2024] adopt more complex strategies such as multi-marginal transport.
Although these multi-view extensions generalize two-view objectives, they still rely solely on pairwise relations and overlook the joint structural relationships among the full set of views.
To overcome these limitations, we introduce a novel similarity function that directly models the joint structure among all $M$ augmented views of an instance. Rather than reducing multi-view interactions to pairwise comparisons or expensive transport computations, our method first estimates a probability distribution over the views on the unit hypersphere and then computes the similarity between two instances via a negative divergence measure between their respective distributions. In particular, we employ the von Mises–Fisher distribution, which is parameterized by a mean direction $\mu$ and a concentration parameter $\kappa$. The mean direction $\mu$ represents the central tendency of the distribution, while the concentration parameter $\kappa$ captures the joint structure among the multi-view features. Moreover, under the special case of two views, our divergence based similarity exactly recovers the cosine similarity used in the InfoNCE loss, thereby bridging the gap between conventional pairwise and multi-view contrastive paradigms.
Our main contributions are:
  • Divergence-based multi-view Contrastive Learning. we propose a Divergence-based Similarity Function(DSF) that captures joint structure across all views without decomposing into pairwise terms.
  • Relation with original InfoNCE loss. We show that, when restricted to two views under specific condition, our similarity recovers cosine similarity in InfoNCE loss, thus offering a unified view of pairwise and multi-view contrastive methods.
  • Hyperparameter-free temperature control We mathematically establish the role of the temperature term in contrastive loss and show that our model learns optimally without manual temperature adjustment.
  • Experimental Results. We validate the effectiveness of our method through extensive experiments on kNN classification, linear evaluation, and resource analysis in terms of GPU memory and training time, consistently outperforming the state-of-the-art contrastive learning approaches.
The remainder of this paper is structured as follows. Section 2 provides an overview of contrastive learning and its multi-view extensions. Section 3 reviews related work on contrastive frameworks with both two-view ($M=2$) and multi-view ($M\geq 3$) settings. In Section 4, we present our proposed divergence-based similarity formulation. Section 5 describes the experimental setup and results. Finally, Section 6 concludes the paper and discusses potential directions for future research.
2 Preliminaries
(a) kNN Top-1 Accuracy with respect to training time.
(b) Linear evaluation accuracy with respect to training time.
Figure 3: More Time Cost Comparison of performance over training time for kNN and linear evaluation. Both figures report accuracy for different training epochs and wall-clock time under a fixed GPU memory budget, with the number of views set to 8 ($M=2m=8$). The right-hand side of each figure shows the elapsed time of MoCo baselines, while the x-axis indicates the time and epochs for the multi-view methods. Note that the lowest dashed line (800 epochs, 5.5h) corresponds to the same GPU memory and training time as the first point(250 epochs, 5.5h) of each multi-view method. Compared to Loss_avg and Feature_avg at 800 epochs, our proposed DSF achieves more than 2$\times$ faster convergence in the kNN setting and over 3$\times$ faster convergence in the linear evaluation setting.
2.1 Contrastive Learning
Contrastive Learning (CL) has emerged as a powerful paradigm in unsupervised and self-supervised learning, aiming to learn representations by distinguishing positive and negative pairs. Here, positives refer to semantically similar pairs whereas negatives to dissimilar ones. The core idea is to map positive samples closer in the embedding space while pushing negative samples apart. Given a positive pair and a set of negative samples, this loss encourages the model to assign high similarity to the positive pair relative to the negatives. The loss can be formulated as follows:
$L_{i}=-\log\frac{e^{(sim(z_{i},z^{+}_{i})/\tau)}}{e^{(sim(z_{i},z^{+}_{i})/\tau)}+\sum\limits_{i\neq j}^{K}e^{(sim(z_{i},z^{-}_{j})/\tau)}}$ (1)
where $z^{+}_{i}$ is the embeddings of the positive pair respect to $z_{i}$ and $z^{-}_{j}$s are embeddings of $K-1$ negative pairs. $sim(\cdot,\cdot$) is a similarity measure (e.g., dot product) and $\tau$ is a temperature parameter.
2.2 von Mises-Fisher Distribution
The von Mises-Fisher (vMF) distribution is a probability distribution dealing with a data that lies on a unit hypersphere. Defined over a $(p-1)$-dimensional unit hypersphere $\mathbb{S}_{p-1}$, vMF distribution describes the probability density of a unit vector with a specified mean direction. The distribution is characterized by a mean direction, indicating the central orientation of the vector, and a concentration parameter, which describes how tightly the distribution clusters around this mean direction.
The probability density function of the vMF distribution for a unit vector $x\in S_{p-1}$ is given by:
$f_{p}(x;,\mu,\kappa)=C_{p}(\kappa)\exp(\kappa\mu^{T}x)$ (2)
where $\mu$ is the mean direction, $\kappa$ is the concentration parameter, and $C_{p}(\kappa)$ is a normalization constant.
A higher concentration parameter $\kappa$ results in samples that are tightly clustered around the mean direction $\mu$, while $\kappa=0$ corresponds to a uniform distribution over the hypersphere.
3 Related Works
3.1 Original Contrastive Learning
Contrastive Learning appears in many self-supervised learning methods. SimCLR[Chen et al., 2020a] leverages the InfoNCE objective[Oord et al., 2018] to pull together representations of two augmented views while pushing apart all other samples. MoCo[He et al., 2020] maintains a momentum encoder, which is a moving average of the parameters to supply a large, consistent set of negatives without labels. [Yeh et al., 2022] decouples the positive and negative terms in the InfoNCE loss, removing negative-positive coupling to enable stable, high performance training with smaller batches and without a momentum encoder. MoCo v3 [Chen et al., 2021] explores contrastive learning using a transformer-based model, specifically the Vision Transformer (ViT) [Dosovitskiy et al., 2020].
3.2 Contrastive Learning with $M\geq 3$ views
Existing $M\geq 3$ multi‐view contrastive learning methods have broadly fallen into two categories. The first category shapes the loss by averaging pairwise terms across views. [Shidani et al., 2024] generate $M$ views per sample and aggregate their pairwise losses either arithmetically or geometrically. The second category is a method that deal with view representations at the feature level before computing the final objective. FastSiam[Pototzky et al., 2022] averages the other $M-1$ view features to define each view’s target , and M3G[Piran et al., 2024] adopts a multi‐marginal optimal‐transport loss by measuring the gap between the true $M$-tuple transport cost and its optimal polymatching cost. In contrast to these approaches, our method integrates multi-views into a single collective loss formulation without incurring additional computational overhead.
4 Divergence-based similarity function
In this section, we explain the details of proposed divergence-based similarity function(DSF) and relations with cosine similarity function. In Sec4.1, we formalize notations throughtout this paper. In Sec4.2, we describe how to estimate vMF distribution parameters $\mu,\kappa$. In Sec4.3,we propose the adjustment to kappa value for computational stabilization. In Sec4.4, we describe the calculation of DSF. In Sec4.5, we show that KL divergence-based similarity function between vMF distribution becomes equivalent to consine similarity function in InfoNCE Loss under certain conditions. Furthermore, we show why our divergence similairty function does not need a temperature hyperparameter and why cosine similarity function need temperature hyperparamter.
4.1 Notation
We denote the batch size, number of views, and feature dimensionality by $B$, $M$, and $D$, respectively. Each encoder is given $m$ views, where $m$ indicates the number of augmented views per encoder, resulting in a total of $M=2m$ views.
The Kullback–Leibler (KL) divergence is written as $D_{\mathrm{KL}}(\cdot\|\cdot)$, and we denote a von Mises–Fisher (vMF) distribution by the calligraphic symbol $\mathcal{D}$.
Let $x$ be an image instance, and $x^{l}$ denote its $l$-th augmented view. Given an encoder $f_{\theta}(\cdot)$, the corresponding feature on the unit hypersphere is computed as $z^{l}=f_{\theta}(x^{l})/\|f_{\theta}(x^{l})\|$, $l=1,\cdots,m$.
contrastive Learning(CL) makes $\mathbb{R}^{B\times 2\times D}$ features with two augmentation with each instance. Multi-view contrastive learning makes $\mathbb{R}^{B\times 2m\times D}$ features with total $M=2m$ views with each instance.
In CL, there are $2B$ features $z_{i}$, $i=1,\cdots,2B$. $z_{i}$ and $z_{B+i}$ are positively correlated. In Multi-view CL, there are $B\cdot 2m$ features $z_{i}^{l}$, $i=1,\cdots,B$ and $l=1,\cdots,m$. $\{z_{i}^{l}\}_{l=1}^{m}$ and $\{z_{B+i}^{l}\}_{l=1}^{m}$ are positively correlated.
Dataset Base Model Method Epoch kNN↑ Linear Evaluation↑
ImageNet-100 MoCo v3 - 300 74.08 81.52
MoCo v3 loss_avg 100 76.62 84.24
MoCo v3 fea_avg 100 76.94 84.06
MoCo v3 DSF 100 79.00 85.66
CIFAR10 MoCo - 800 88.44 88.10
MoCo loss_avg 250 88.28 88.58
MoCo fea_avg 250 87.47 88.02
MoCo DSF 250 90.04 91.21
Table 1: Comparison of kNN and Linear Evaluation Results Under Equivalent Resource Settings. All experiments are conducted under the same GPU memory and time constraints. The number of views is fixed at $M=2m=8$.
4.2 Estimate vMF Distribution
Before introducing the DSF, we first describe the process of constructing von Mises–Fisher (vMF) distributions from multiple augmented views of a single image instance.
Given an image instance $x$, we generate $M$ augmented views via data augmentation. These views are divided into two groups of $m$ samples each. All augmented images are passed through the encoder to obtain $M$ feature vectors on the unit hypersphere. Each group of features is then used to estimate the parameters of a vMF distribution. As a result, each instance is represented by two vMF distributions, each constructed from one group of augmented features.
Let $z^{1},\dots,z^{m}$ denote the feature vectors corresponding to the $m$ augmented views of an instance $x$. The mean direction $\mu$ and concentration parameter $\kappa$ of the vMF distribution are estimated as follows:
$\mu=\frac{\bar{z}}{\bar{R}},\quad\kappa=A_{p}^{-1}(\bar{R}),$ (3)
where
$\bar{z}=\frac{1}{m}\sum\limits_{l=1}^{m}z^{l},\quad\bar{R}=\|\bar{z}\|.$ (4)
Since computing $\kappa=A_{p}^{-1}(\bar{R})$ is intractable in practice, we instead adopt an approximate form of $\kappa$ proposed by [Banerjee et al., 2005]:
$\kappa=\frac{\bar{R}(p-\bar{R}^{2})}{1-\bar{R}^{2}}$ (5)
Consequently, the vMF distribution constructed from the features $z^{1},\dots,z^{m}$ is denoted as:
$\mathcal{D}(z^{1},\dots,z^{m})=\mathcal{D}(\mu,\kappa).$ (6)
For subscript indexing $i$ :
$\mathcal{D}(z_{i}^{1},\dots,z_{i}^{m})=\mathcal{D}(\mu_{i},\kappa_{i})=\mathcal{D}_{i}.$ (7)
4.3 Stabilization of the Concentration Parameter $\kappa$
As a final step before introducing the DSF, we describe a strategy to stabilize the concentration parameter $\kappa$, which is essential for maintaining numerical stability during training.
As training progresses, the value of $\kappa$ can grow excessively, often leading to numerical overflow in computations. This issue arises from two primary factors:
  • High feature dimensionality. For a fixed mean resultant length $\bar{R}$, a larger feature dimension $p$ results in a sharper increase in $\kappa$. This becomes problematic, particularly when using higher-dimensional feature spaces (e.g., $D=128$ in our setup), as commonly done in larger models.
  • Feature concentration. As the encoder learns more discriminative representations, the extracted features become increasingly concentrated, causing $\bar{R}$ to approach 1, which in turn further amplifies $\kappa$.
Although this behavior is theoretically valid, it leads to numerical instability when computing divergence measures in practice. To mitigate this issue, we apply two simple yet effective techniques:
  1. 1.
    Dimension-based normalization: We normalize $\kappa$ by the feature dimension $p$, as $\kappa$ grows approximately linearly with $p$ when the mean resultant length $\bar{R}$ is fixed, as shown in Eq. 5. This prevents dimension-driven growth.
  2. 2.
    Scaling of mean resultant length: To avoid $\bar{R}$ approaching 1, we apply a scaling factor $\lambda_{\bar{R}}\in(0,1)$. In our experiments, we use $\lambda_{\bar{R}}=0.95$, modifying $\bar{R}$ as follows:
    $\bar{R}\longrightarrow\lambda_{\bar{R}}\cdot\bar{R}$ (8)
These modifications are empirically shown to stabilize training and avoid divergence explosion in high-dimensional settings. Further details and illustrative figures can be found in the Appendix C Figure 6.
4.4 Divergence-based similarity function
4.4.1 Similarity function
Similarity function $\text{sim}(\cdot,\cdot)$ is given by $\text{sim}(\cdot,\cdot):\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$. For example, cosine similarity function is given as $\text{sim}_{\text{cos}}(\cdot,\cdot):\mathbb{S}_{p-1}\times\mathbb{S}_{p-1}\rightarrow\mathbb{R}$, $\text{sim}_{\text{cos}}(x,y)=x^{T}y$. Accordingly, DSF handles multi-view features, and its similarity function is defined as $\text{sim}_{\text{div}}(\cdot,\cdot):\mathbb{S}_{p-1}^{m}\times\mathbb{S}_{p-1}^{m}\rightarrow\mathbb{R}$.
4.4.2 KL Divergence of vMF Distribution
Let $\mathcal{D}_{i}=\mathcal{D}(\mu_{i},\kappa_{i})$ and $\mathcal{D}_{j}=\mathcal{D}(\mu_{j},\kappa_{j})$ be any two vMF distribution. The KL divergence value of $\mathcal{D}_{i}$ and $\mathcal{D}_{j}$ is expressed as the sum of three terms.
$\begin{split}&D_{\mathrm{KL}}(\mathcal{D}_{i}||\mathcal{D}_{j})\\ &=\log\frac{f_{p}(A_{p}(\kappa_{i})\mu_{i};\mu_{i},\kappa_{i})}{f_{p}(A_{p}(\kappa_{i})\mu_{i};\mu_{j},\kappa_{j})}\\ &=\left(\frac{p}{2}-1\right)\log\frac{\kappa_{i}}{\kappa_{j}}+\log\frac{I_{\frac{p}{2}-1}(\kappa_{j})}{I_{\frac{p}{2}-1}(\kappa_{i})}\\ &\qquad\qquad\qquad+A_{p}(\kappa_{i})\cdot\left(\kappa_{i}-\kappa_{j}\cdot\mu_{i}^{T}\mu_{j}\right)\\ \\ \end{split}$ (9)
The KL divergence range in from 0 to infinity. KL divergence is not a distance function of two distribution. As the KL divergence value close to zero, it means the distributions are similar i.e. $\mu,\kappa$ become close. So, our proposed divergence-based similarity function is negative KL divergence value as following.
$\text{sim}_{\text{div}}(\mathcal{D}_{i},\mathcal{D}_{j})=-D_{\mathrm{KL}}(\mathcal{D}_{i}||\mathcal{D}_{j})$ (10)
$L_{i,\text{InfoNCE}}^{\text{div}}=-\log\left(\frac{e^{\text{sim}_{\text{div}}(\mathcal{D}_{i},\mathcal{D}_{i}^{+})}}{e^{\text{sim}_{\text{div}}(\mathcal{D}_{i},\mathcal{D}_{i}^{+})}+\sum\limits_{j\neq i}^{K}e^{\text{sim}_{\text{div}}(\mathcal{D}_{i},\mathcal{D}_{j})}}\right)$ (11)
Here, $\mathcal{D}_{i}^{+}$ denotes the positive distribution paired with $\mathcal{D}_{i}$. The negative distributions are denoted by $\mathcal{D}_{j}$, where $K$ represents the number of negative distributions. Note that the value of $K$ may vary depending on the training framework.
4.5 Analysis : Relation with Cosine Similarity Function
Next, we establish the connection between the conventional cosine similarity function and our proposed DSF. We further explain why cosine similarity requires a temperature scaling term for effective optimization, whereas our DSF inherently eliminates the need for such adjustment.
Theorem 1 (Relation between Divergence-based and Cosine Similarities).
Let $Z_{i}=\{z^{1}_{i},\cdots,z^{m}_{i}\}$ and $Z_{j}=\{z^{1}_{j},\cdots,z^{m}_{j}\}$ are $m$ multi-view features. $\mathcal{D}_{i}$ and $\mathcal{D}_{j}$ are the estimated von Mises-Fisher distribution from $Z_{i}$ and $Z_{j}$. Let $\text{sim}_{\text{cos}}(\cdot,\cdot)$ is cosine similarity and $\text{sim}_{\text{div}}(\cdot,\cdot)$ is the negative Kullback-Leibler divergence between two vMF distribution. If m=1, $\kappa_{1}$=$\kappa_{2}$ and $A_{p}(\kappa_{1})\cdot\kappa_{1}=1/\tau$ then
$\text{sim}_{\text{div}}(\mathcal{D}(z_{i}),\mathcal{D}(z_{j}))=\text{sim}_{\text{cos}}(z_{i},z_{j})/\tau+constant$ (12)
Furthermore, InfoNCE Loss with $sim_{\text{div}}$ becomes equivalent to that using $sim_{\text{cos}}$ one:
$L^{\text{div}}_{\text{InfoNCE}}=L^{\text{cos}}_{\text{InfoNCE}}$ (13)
Proof.
Proof is at Appendix A. ∎
Theorem 1 says that when we have two views and kappa is set under certain conditions. Then our proposed divergence-based simiarity function becomes cosine simiarity function with some constant value. Furthermore, this leads to fact that InfoNCE loss becomes equivalent.
Now, we explain the dependency of cosine similarity on temperature hyperparameter and independence of our DSF with respect to temperature hyperparameter.
Proposition 1.
Let $K$ be the number of negative pairs, and assume all negative similarities have the same value. Let $s^{+}$ and $s^{-}$ denote the positive and negative similarities, respectively. Then, the InfoNCE loss becomes:
$\begin{split}L_{\text{InfoNCE}}&=-\log\frac{\exp(s^{+}-s^{-})}{\exp(s^{+}-s^{-})+K}\\ (s^{+}-s^{-})&\in\left\{\begin{matrix}[-\frac{2}{\tau},\frac{2}{\tau}]&\text{if }\text{sim}=\text{sim}_{\text{cos}}\\ (-\infty,\infty)&\text{if }\text{sim}=\text{sim}_{\text{div}}\end{matrix}\right.\end{split}$ (14)
In the optimal case for cosine similarity, we have $s^{+}=+1$ and $s^{-}=-1$, resulting in:
$\begin{split}L_{\text{InfoNCE}}&=-\log\frac{\exp(2/\tau)}{\exp(2/\tau)+K}\\ &=6.3196,\quad\tau=1.0,\,K=4096\\ &=4.3310,\quad\tau=0.5,\,K=4096\\ &=0.1706,\quad\tau=0.2,\,K=4096\\ &=0.0000,\quad\tau=0.1,\,K=4096\\ \end{split}$ (15)
It has been theoretically shown (see Proposition1) that cosine similarity cannot represent a appropriate loss value without the use of a temperature term $\tau$. The core issue arises from the limited range of the similarity margin $(s^{+}-s^{-})$, which restricts the contrast between positive and negative pairs. To compensate for this limitation, the temperature parameter $\tau$ is introduced to amplify this margin, allowing for better gradient signals during training.
The number of negative samples $K$ varies depending on the contrastive learning framework. For instance, in SimCLR and DCL, $K=2B-2=254$ when the batch size $B=128$, while in MoCo, $K$ is determined by the size of the memory queue, ranging from 4,096 to 65,536 depending on the dataset.
As summarized in Table 2, the InfoNCE loss yields inappropriate values without the temperature parameter $\tau$ in all $K$ values. It says cosine similarity requires a carefully tuned temperature parameter across all frameworks to function effectively. In contrast, our proposed DSF expresses an appropriate loss without the need for such a temperature hyperparameter. See Figure 5 for $(s^{+}-s^{-})$ margin in our DSF experiment.
$L_{\text{InfoNCE}}$
$\tau$ K=256 K=4096 K=65536
1.0 3.573 6.319 9.090
0.5 1.738 4.331 7.091
0.2 0.011 0.170 1.380
0.1 0.000 0.000 0.0001
Table 2: InfoNCE loss values under optimal similarity ($s^{+}=1$, $s^{-}=-1$) for varying temperature $\tau$ and the number of negative pairs $K$.
Algorithm 1 Three Multi-View Methods
Input: Instance $x_{i}\in\mathbb{R}^{C\times H\times W}$, encoders $f_{\theta}$ and $f_{\theta^{\prime}}$.
Parameter: Number of views $M=2m$.
Output: Divergence-based similarity between positive pairs.
1:  Apply $M=2m$ augmentations to $x_{i}$, yielding $x_{i}^{1},\dots,x_{i}^{m}$ and $x_{B+i}^{1},\dots,x_{B+i}^{m}$
2:  Extract features:
$z_{i}^{l}=\frac{f_{\theta}(x_{i}^{l})}{\|f_{\theta}(x_{i}^{l})\|},\quad z_{B+i}^{l}=\frac{f_{\theta^{\prime}}(x_{B+i}^{l})}{\|f_{\theta^{\prime}}(x_{B+i}^{l})\|}\quad$
for $\text{for }l=1,\dots,m$
3:  if multi-view method = loss_avg:
$L_{\text{loss\_avg}}=\frac{1}{m^{2}}\sum_{l=1}^{m}\sum_{l^{\prime}=1}^{m}L_{\text{InfoNCE}}^{\cos}(z_{i}^{l},z_{B+i}^{l^{\prime}})$
4:  if multi-view method = fea_avg:
$z_{i}^{\text{avg}}=\frac{1}{m}\sum_{l=1}^{m}z_{i}^{l},\quad z_{B+i}^{\text{avg}}=\frac{1}{m}\sum_{l=1}^{m}z_{B+i}^{l}$
$L_{\text{fea\_avg}}=L_{\text{InfoNCE}}^{\cos}(z_{i}^{\text{avg}},z_{B+i}^{\text{avg}})$
5:  if multi-view method = DSF:  Estimate vMF distribution parameters as in Eq. (3)–(4):
$(\mu_{i},\kappa_{i})\leftarrow\{z_{i}^{1},\dots,z_{i}^{m}\}$
$(\mu_{B+i},\kappa_{B+i})\leftarrow\{z_{B+i}^{1},\dots,z_{B+i}^{m}\}$
6:   Compute divergence-based similarity:
$\mathrm{sim}_{\mathrm{div}}(\mathcal{D}_{i},\mathcal{D}_{B+i})=-D_{\mathrm{KL}}(\mathcal{D}_{i}\,\|\,\mathcal{D}_{B+i})$
$L_{\text{DSF}}=L_{\text{InfoNCE}}^{\mathrm{div}}\left(\mathcal{D}(\mu_{i},\kappa_{i}),\mathcal{D}(\mu_{B+i},\kappa_{B+i})\right)$
5 Experiments
5.1 Experimental Setup
Every experiment is conducted with one or four NVIDIA GeForce RTX 3090.
Dataset
This study utilizes two benchmark datasets: CIFAR-10 and ImageNet-100. CIFAR-10 [Krizhevsky et al., 2009] consists of 60,000 color images of size $32\times 32$ pixels, divided into 50,000 training and 10,000 test images across 10 object classes. ImageNet-100 is a curated subset of the ImageNet dataset, comprising 100 randomly selected classes with approximately 1,300 images per class. It serves as a mid-scale benchmark for evaluating representation learning performance.
These datasets are used to validate the proposed method on both small-scale and mid-scale visual recognition tasks.
Model Arichitecture
We primarily follow the MoCo v2 [Chen et al., 2020b] framework for contrastive learning. Throughout this paper, any reference to ”MoCo” refers to MoCo v2 unless otherwise specified. We adopt ResNet-18 as the encoder for experiments on CIFAR-10 The final fully connected layer projects features into a 128-dimensional embedding space.
For large-scale evaluation, we extend our experiments to ImageNet-100 using MoCo v3 [Chen et al., 2021] with a Vision Transformer (ViT-S/16) backbone. This enables assessment of the scalability of the proposed method across both convolutional and transformer-based architectures.
Data Augmentation
For data augmentation, we apply a combination of RandAugment [Cubuk et al., 2020], ColorJitter, Random Grayscale, Gaussian Blur, and Horizontal Flip, following common practice in contrastive learning setups.
5.2 Evaluations
To evaluate the model’s performance, we conducted following tasks: k-Nearest Neighbors (kNN) and Linear Evaluation. Linear Evaluation assesses the representational quality of the encoder by attaching a classifier, while kNN measures similarity-based performance directly within the encoder’s embedding space.
kNN Evaluation.
We evaluated the model with a kNN approach to measure the encoder’s capacity to retain similarity in the embedding space. Evaluation was conducted on CIFAR-10, ImageNet-100 using the output feature of the trained query encoder. The K value was set to 200, enabling an assessment of the encoder’s representations in a non-parametric, unsupervised manner. In our experiments, the proposed model achieved a kNN accuracy of 79.00% on the ImageNet-100 dataset, outperforming the baseline MoCo v3 model, which attained 74.08%. On the CIFAR-10 dataset, our model achieved 90.04% accuracy, compared to 88.44% with the MoCo baseline.
Linear Evaluation.
To further validate the effectiveness of our method, we evaluate the quality of the learned representations through linear evaluation. Specifically, after performing self-supervised pretraining on ImageNet-100 and CIFAR-10 using our proposed DSF framework, we freeze the encoder and train a linear classifier (a single-layer MLP) on top of the extracted features.
On the ImageNet-100 dataset, our method achieved a linear evaluation accuracy of 85.66%, outperforming the baseline MoCo v3 model, which attained 81.52%. Similarly, on the CIFAR-10 dataset, our method achieved an accuracy of 91.21%, compared to 88.10% obtained by the MoCo baseline. These results demonstrate that DSF leads to stronger and more transferable representations across different datasets.
(a) kNN Top-1 Accuracy with respect to number of view.
(b) Linear Evaluation Top-1 Accuracy with respect to number of view.
Figure 4: Comparison of kNN and linear evaluation accuracies according to the number of views in three multi-view learning methods.
5.3 Same Resource Comparison
Our method samples $M$ multiple augmented views from each image, which could potentially increase computational overhead. Therefore, it is essential to ensure that the proposed approach remains efficient under equivalent computational constraints.
Computational Efficiency.
Our method samples $M$ augmented views from each image instance, which could potentially increase computational overhead. To address this, we ensure fair comparisons by maintaining consistent computational budgets across all experiments.
To ensure the practicality of our method, we carefully control and match computational budgets across all experiments. Specifically, we fix the product of batch size and number of views—i.e., $B\times M$—so that memory usage remains constant regardless of the number of views. For example, in our ImageNet-100 experiments (see Table 1), the baseline MoCo is trained with a configuration of $512\times 2$, whereas our method and other multi-view baselines use $128\times 8$, maintaining identical GPU memory usage. This configuration strategy is also consistently applied to CIFAR-10 experiments.
Moreover, to equalize the overall training time, we proportionally reduce the number of training epochs for multi-view methods. As a result, the total runtime difference between our approach and baseline models is strictly limited to within 30 minutes—less than 4% and 10% of the entire training duration in the ImageNet-100 and CIFAR-10 setting. These measures ensure that the observed performance improvements are attributable to our proposed DSF method itself, rather than to increased computational resources.
5.4 More Resource Comparison
5.4.1 More Time Cost: Faster Convergence with Divergence Similarity
Self-supervised learning methods are known to require longer training times for convergence due to the nature of instance discrimination objectives and the inherent challenge of learning meaningful representations from unlabeled data. This issue becomes even more pronounced when dealing with complex datasets or limited batch sizes.
To investigate convergence behavior, we extended the number of training epochs and compared model accuracy as a function of elapsed training time, as shown in Figure 3. While the baseline MoCo gradually improves over a long training schedule (55 hours at 8000 epochs), our proposed DSF reaches comparable accuracy within only 11 hours at 400 epochs, using merely 20% of the total training time. Specifically, compared to the Loss_avg and Feature_avg baselines trained for 800 epochs, our DSF reaches similar accuracy within just 400 epochs, achieving a 2× speedup in convergence. Moreover, we observe that the performance gap between our DSF and the two other multi-view baselines (Loss_avg, Feature_avg) remains consistently large, regardless of the total training time. This indicates that our method not only converges faster but also achieves a higher final accuracy that is not easily matched even with extended training.
This result highlights that our divergence-based formulation facilitates more efficient optimization in self-supervised learning. Notably, all three multi-view methods considered in this comparison have identical per-epoch training time, ensuring a fair evaluation.
5.4.2 More GPU Memory : Number of Views
We sample multiple augmented views from each image; thus, it is important to investigate how the number of sampled views affects the quality of the learned representations. Prior studies [Tian et al., 2020, Shidani et al., 2024, Pototzky et al., 2022] have shown that increasing the number of augmentations generally leads to better representation quality.
In our framework, the number of views directly impacts the accuracy of the estimated von Mises–Fisher (vMF) distribution. As the number of views increases, the estimated parameters $\mu$ and $\kappa$ become more reliable due to reduced estimation variance.
Figure 4 presents the performance trends of three multi-view learning methods on both kNN and linear evaluation tasks, as the number of views $M$ increases. Across all view counts, DSF consistently outperforms the other two approaches. These results suggest that our divergence-based formulation provides stable and reliable improvements in multi-view learning settings.
5.5 Temperature Free
In numerous studies, the temperature parameter $\tau$ has been shown to be critical for achieving optimal performance in contrastive learning [Zhang et al., 2021, Zhang et al., 2022, Kukleva et al., 2023]. In our method, however, $\tau$ is not required, as discussed in Proposition 1, and it can even degrade performance when introduced.
To validate this, we conducted experiments over a wide range of temperature values, from 0.1 to 5.0. Interestingly, setting $\tau=1$ (i.e., using the raw similarity without scaling) consistently resulted in the best performance in Table 3. This suggests that our DSF is inherently numerically stable and does not require additional temperature tuning.
Temperature $\tau$ 0.8 0.9 1 1.1
kNN 91.2 91.3 91.8 91.4
Linear Evaluation 92.0 92.3 92.5 92.0
Table 3: Temperature Experiment. MoCo + DSF results in CIFAR10.
5.6 Training Details
To train on ImageNet-100 dataset, MoCo v3 with ViT model as the encoder from scratch. We use AdamW as our optimizer. The weight decay is 0.1 and the encoder momentum is 0.99 for updating encoder. We choose batch size of 512, learning rate of 0.03. In our proposed method, a larger learning rate of 0.06 is used. GPU used for overall training is NVIDIA GeForce RTX 3090.
6 Future Work
Figure 5: Training dynamics of the similarity margin $(s^{+}-s^{-})$. Values are averaged over mini-batches throughout training. This plot shows the similarity margin of MoCo v3 with DSF ($128\times 8$) configuration on ImageNet-100.
We use von Mises-Fisher distribution and Kullback-Leibler divergence for our proposed divergence-based similarity function. Although they are popular choice, there are plenty of available distributions and divergences. For example, in divergence, there are other divergences such as Rényi divergence, $\chi^{\alpha}$divergence, and Jensen–Shannon divergence—collectively known as the family of $f$-divergences could be explored. Investigating how these alternative divergences influence the representation quality and computational efficiency of contrastive methods remains a promising direction for future work.
Acknowledgements
This work is supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (RS-2024-00421203, RS-2024-00406127) and MSIT/IITP(2021-0-00077)
Appendix A Appendix: Proof of Theorem 1
We prove that the divergence-based similarity function becomes equivalent to the cosine similarity function in InfoNCE loss under specific conditions.
Assume $m=1$, $\kappa_{1}=\kappa_{2}=\kappa$, and $A_{p}(\kappa)\cdot\kappa=1/\tau$. Then, the divergence-based similarity simplifies as:
$\text{sim}_{\text{div}}(\mathcal{D}(z_{i}),\mathcal{D}(z_{j}))=\frac{1}{\tau}\text{sim}_{\text{cos}}(z_{i},z_{j})+C,$ (16)
where $C$ is a constant independent of $(z_{i},z_{j})$ pair.
Substituting this into the InfoNCE loss:
$\begin{split}&L_{i,\text{InfoNCE}}^{\text{div}}\\ &=-\log\left(\frac{e^{\text{sim}_{\text{div}}(\mathcal{D}_{i},\mathcal{D}_{i}^{+})}}{e^{\text{sim}_{\text{div}}(\mathcal{D}_{i},\mathcal{D}_{i}^{+})}+\sum\limits_{j\neq i}^{K}e^{\text{sim}_{\text{div}}(\mathcal{D}_{i},\mathcal{D}_{j})}}\right)\\ &=-\log\left(\frac{e^{\frac{1}{\tau}\text{sim}_{\text{cos}}(z_{i},z_{i}^{+})+C}}{e^{\frac{1}{\tau}\text{sim}_{\text{cos}}(z_{i},z_{i}^{+})+C}+\sum\limits_{j\neq i}^{K}e^{\frac{1}{\tau}\text{sim}_{\text{cos}}(z_{i},z_{j})+C}}\right)\\ &=-\log\left(\frac{e^{\frac{1}{\tau}\text{sim}_{\text{cos}}(z_{i},z_{i}^{+})}}{e^{\frac{1}{\tau}\text{sim}_{\text{cos}}(z_{i},z_{i}^{+})}+\sum\limits_{j\neq i}^{K}e^{\frac{1}{\tau}\text{sim}_{\text{cos}}(z_{i},z_{j})}}\right)\\ &=L_{i,\text{InfoNCE}}^{\text{cos}}.\end{split}$ (17)
Appendix B Feature Averaging Also Encodes Only Pairwise Relationships
Loss_avg is clearly a pairwise-based method. We further demonstrate that Feature_avg, despite aggregating multiple views, still considers only pairwise relationships among features, as can be shown through a straightforward calculation.
$z_{i}^{\text{avg}}$ $=\frac{1}{m}\sum_{l=1}^{m}z_{i}^{l},\quad z_{B+i}^{\text{avg}}=\frac{1}{m}\sum_{l=1}^{m}z_{B+i}^{l}$
$\text{sim}_{\text{cos}}(z_{i}^{\text{avg}},z_{B+i}^{\text{avg}})$ $=(z_{i}^{\text{avg}})^{T}z_{B+i}^{\text{avg}}$
$=\left(\frac{1}{m}\sum_{l=1}^{m}z_{i}^{l}\right)^{T}\left(\frac{1}{m}\sum_{l^{\prime}=1}^{m}z_{B+i}^{l^{\prime}}\right)$
$=\frac{1}{m^{2}}\sum_{l=1}^{m}\sum_{l^{\prime}=1}^{m}(z_{i}^{l})^{T}z_{B+i}^{l^{\prime}}$ (18)
As shown above, the final similarity score is an unweighted average over all pairwise cosine similarities between individual view features $z_{i}^{l}$ and $z_{B+i}^{l^{\prime}}$. Thus, Feature_avg does not model joint structure beyond pairwise relationships.
Appendix C Stabilization of the Concentration Parameter $\kappa$
To address numerical instability during training, we use the approximated concentration parameter $\kappa$ as defined in Equation 5. However, even this approximation can lead to excessively large values of $\kappa$ as training progresses. To mitigate this, we propose two stabilization techniques.
Figure 6 illustrates the original approximated $\kappa$ values computed from Equation 5, alongside the stabilized values obtained after applying our proposed techniques.
Figure 6: Consistency by Kappa Normalization. $\kappa$ value is calculated as approximated formulation as (5). $x$-axis is $\bar{R}$ and $y$-axis is calculated $\kappa$. $p$ is feature dimension size. In our backbone ResNet18, feature dimension is 128. (upper) Directly calculated $\kappa$. (lower) $\kappa$ is divided by feature dimension $p$ and calculated with $\bar{R}$ multiplied by $\lambda_{\bar{R}}$. Without these two techniques, $\kappa$ value is going excessively large.